If x has a normal distribution with a mean 15 and standard deviation 3, describe
_
the distribution of x values for sample size n, where n=4, n=16, and n =100. How
_
do the x distributions compare for the various sample sizes?

Answer by stanbon(75887)

If x has a normal distribution with a mean 15 and standard deviation 3, describe the distribution of x values for sample size n, where n=4, n=16, and n =100. How do the x distributions compare for the various sample sizes?
----
The mean of the sample means of any size = the mean of the population.
---
The std of the sample means for samples of size n = s/sqrt(n) where
s is the std of the population.
------------------------------------
mean for all your distributions is 15
---
std for samples of size 4 is 3/sqrt(4) = 3/2
---
std of samples of size 16 is 3/sqrt(16) = 3/4
---
std of samples of size 100 = 3/sqrt(100) = 3/10
-----
Notice: As the samples size gets larger the spread of the
distribution gets smaller, or tighter.
=======================================
Cheers,
Stan H.
===========